OpenStudy (anonymous):

Which of the following is a polynomial with roots -3, -3i, and 3i ? x3 + 3x2 + 9x + 27 x3 - 3x2 + 9x - 27 x3 + 9x2 + 9x + 81 x3 - 9x2 + 9x - 81

OpenStudy (triciaal):

I got the first one (x-3i)(x + 3i) = x^2 - 9i^2 = x^2 + 9 (x^2 + 9 )(x + 3) = x^3 + 9x + 3 x^2 + 27 = x^3 + 3 x ^2 + 9 x + 27 = option 1

OpenStudy (anonymous):

Ok @JordanGonen First let's ask, what do the factors of -3, -3i, and 3i look like? If I said the factors of -2 and 4 were (x+2) and (x-4) What would be your best guess here for the -3, the -3i, and 3i?

OpenStudy (anonymous):

-3 = (x+3) But I have no clue when it comes to the imaginary numbers. Unless its (x+-3i), which highly doubt.

OpenStudy (anonymous):

Right and what is (x+-1) for example?

OpenStudy (anonymous):

(x-1) right because we multiply the signs (+) * (-) and this will always be (-) right?

OpenStudy (anonymous):

So (x+-3i) is what?

OpenStudy (anonymous):

(x-3i)?

OpenStudy (anonymous):

It's (x-3i), correct!

OpenStudy (anonymous):

so the factor of -3, -3i, and 3i are?

OpenStudy (anonymous):

Factors*

OpenStudy (anonymous):

(x+3),(x-3i),(x+3i)

OpenStudy (anonymous):

Now let's go over a few things about the imaginary number (i) Great job on showing those factors, those are exactly right.

OpenStudy (anonymous):

Do you know what i is equal too?

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):



OpenStudy (anonymous):

We call it imaginary because we cannot take the square root of a negative number.

OpenStudy (anonymous):

But (i) is always equal to the sqrt(-1) ok?

OpenStudy (anonymous):

Now let me ask this question really quick, do you know how to do this?